# Spinor bilinears and Killing-Yano forms in generalized geometry

### v1

@@ IV Generalized Killing and twistor spinors
where ϵ=±italic-$\epsilon=\pm$ on subbundles $V_{\pm}$. In fact, the second term on the right hand side corresponds to the Clifford bracket action of $-\frac{1}{2}i_{X_{a}}H$ to 1-forms $\widetilde{Y}$ and $\eta$
@@
$[\alpha,\omega]_{Cl}=-2i_{X_{a}}\alpha\wedge i_{X^{a}}\omega$ (35)
and for the cases of 2-form $-\frac{1}{2}i_{X_{a}}H$ and 1-forms $\widetilde{Y}$ and $\eta$, we have
@@
$[i_{X_{a}}H,\mathcal{Y}]_{Cl}=\left(\begin{array}[]{cc}\widetilde{[i_{X_{a}}H,\widetilde{Y}]}_{Cl}\\ {[i_{X_{a}}H,{\eta}]}_{Cl}\end{array}\right).$ (38)
Let us consider a generalized spinor $\psi$ corresponding to an inhomogeneous differential form which satisfies the Killing spinor equation in terms of the generalized connection and generalized 1-forms
@@
where we have used the identities $d^{H}=e^{a}\wedge\nabla^{H}_{X_{a}}$ and $e^{a}\wedge i_{X_{a}}\psi_{p}=p\psi_{p}$. Similarly, by applying $-i_{X^{a}}$ to (48), we obtain
@@
which corresponds to the conformal Killing-Yano (CKY) equation written in terms of a connection modified by a 3-form flux $H$ corresponding to torsion [29, 30, 31]. Here, $H$ corresponds to the skew-symmetric torsion and torsion 2-forms $T^{a}$ are defined as $T^{a}=i_{X^{a}}H$. Then, a solution of the generalized Killing spinor equation corresponds to a solution of the CKY equation with torsion $H$.
Similarly, let us consider the generalized twistor spinor equation
@@
and the generalized Dirac operator acting on generalized spinor $\psi$ corresponds to
@@
$=$ $(e_{a}\wedge+\epsilon i_{X_{a}})\nabla^{H}_{X_{a}}\psi$
@@
So, a solution of the generalized twistor equation corresponds to a solution of the equation (57). In particular, if the generalized twistor spinor $\psi$ consists of a homogeneous degree differential form, namely it only includes the $p$-form component $\psi=\psi_{p}$, then the above equation reduces to
@@ A $p$-form $\omega$ is called a KY $p$-form, if it satisfies the following equation
because of $\mathcal{G}(\mathcal{Y},\mathcal{Z})$ is a function. Since $\mathbb{D}$ is metric compatible $\mathbb{D}_{\mathcal{X}}\mathcal{G}=0$ and from (62) we have
@@
By using torsion-free condition, we can write the Courant bracket $[\mathcal{X},\mathcal{Y}]_{C}$ in terms of the generalized connection $\mathbb{D}^{H}$ and we have
@@
$<d<\mathcal{X},\mathcal{Y}>,\mathcal{Z}>=i_{\mathcal{Z}}d<\mathcal{X},\mathcal{Y}>=\mathbb{D}^{H}_{\mathcal{Z}}<\mathcal{X},\mathcal{Y}>=<\mathbb{D}^{H}_{\mathcal{Z}}\mathcal{X},\mathcal{Y}>+<\mathcal{X},\mathbb{D}^{H}_{\mathcal{Z}}\mathcal{Y}>$ (75)
$<\mathcal{Y},d<\mathcal{X},\mathcal{Z}>>=i_{\mathcal{Y}}d<\mathcal{X},\mathcal{Z}>=\mathbb{D}^{H}_{\mathcal{Y}}<\mathcal{X},\mathcal{Z}>=<\mathbb{D}^{H}_{\mathcal{Y}}\mathcal{X},\mathcal{Z}>+<\mathcal{X},\mathbb{D}^{H}_{\mathcal{Y}}\mathcal{Z}>.$ (76)
@@
$i_{\mathcal{X}}d^{H}\widetilde{\mathcal{Y}}$ $=$ $i_{\mathcal{X}}(\mathcal{E}^{A}\wedge\mathbb{D}^{H}_{\mathcal{X}_{A}}\widetilde{\mathcal{Y}})$ (79)
@@
$\mathbb{D}^{H}_{\mathcal{X}}\widetilde{\mathcal{Y}}$ $=$ $\frac{1}{2}i_{\mathcal{X}}d^{H}\widetilde{\mathcal{Y}}+\frac{1}{2}\mathcal{E}^{A}\wedge\bigg{(}i_{\mathcal{X}}\mathbb{D}^{H}_{\mathcal{X}_{A}}\widetilde{\mathcal{Y}}+i_{\mathcal{X}_{A}}\mathbb{D}^{H}_{\mathcal{X}}\widetilde{\mathcal{Y}}\bigg{)}$ (80)
@@
$-(\mathbb{L}_{\mathcal{Y}}\mathcal{G})(\mathcal{X},\mathcal{X}_{A})=\frac{1}{2}\bigg{(}\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}}\mathcal{Y},\mathcal{X}_{A})+\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}_{A}}\mathcal{Y},\mathcal{X})\bigg{)}+\frac{1}{2}\bigg{(}\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}_{A}}\mathcal{X},\mathcal{Y})+\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}}\mathcal{X}_{A},\mathcal{Y})\bigg{)}.$ (82)
@@
Let us choose $\mathcal{X}=\mathcal{X}_{B}$ as a basis vector, and consider the admissible metric with basis $\mathcal{X}_{A}=\{X_{a}+\epsilon e^{a}\}$ and $\mathcal{E}^{A}=\{e^{a}+\epsilon X_{a}\}$. Then, we have from (20)
@@
with respect to an admissible basis $\mathcal{X}_{A}=\{X_{a}+\epsilon e^{a}\}$. So, (86) corresponds to the generalized Killing equation in terms of the generalized connection. As the anti-symmetric generalization of generalized Killing equation, we can write the generalized KY equation for generalized $p$-forms $\mathcal{\alpha}$ as
@@
where $d^{H}=\mathcal{E}^{A}\wedge\mathbb{D}^{H}_{\mathcal{X}_{A}}=d^{G}+\frac{\epsilon}{4}H$ where $d^{G}$ is defined as $d^{G}=\mathcal{E}^{A}\wedge\nabla_{\pi({\mathcal{X}_{A}})}$. By using (34), (LABEL:eq83) is written explicitly as
$\mathbb{D}^{H}_{\mathcal{X}_{A}}\mathcal{\alpha}=\nabla_{X_{a}}\mathcal{\alpha}-\epsilon\frac{1}{12}[i_{X_{a}}H,\mathcal{\alpha}]_{Cl}=\frac{1}{p+1}i_{X_{a}}\left(d^{G}\mathcal{\alpha}+\epsilon\frac{1}{4}H\wedge\mathcal{\alpha}\right)$ (88)
in terms of flux terms which resembles the KY equation modified by 3-form flux $H$ corresponding to torsion.
VI Relation Between Generalized KY Forms and Generalized Killing Spinors
We will investigate the relation between generalized Killing spinors satisfying (39) and generalized KY forms satisfying (87). If we have two generalized Killing spinors $\phi$ and $\psi$ satisfying (39), we can construct generalized $p$-form bilinears $(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}$ where a $p$-form bilinear is written in terms of the generalized form basis as
@@ VI Relation Between Generalized KY Forms and Generalized Killing Spinors
from the generalized Killing spinor equation. Here, we use the property $\overline{\mathbb{D}_{\mathcal{X}_{A}}\psi}=\mathbb{D}_{\mathcal{X}_{A}}\overline{\psi}$. By using the definition $\overline{\psi}\phi=(\psi,\phi)$, we can write for a generalized spinor $\kappa$
@@
and since $j$ is identity and $\mathcal{J}=\xi\eta$ for generalized spinors, we have
@@
$d^{G}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ $2\lambda(p+1)\left[(\psi\overline{\phi})_{p+1}+(\phi\overline{\psi})_{p+1}\right]$ (102)
$\delta^{G}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ 00 0 (103)
@@
which correspond to the generalized KY equation. For $p$ is even, we have
@@
$d^{G}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ 00 0 (106)
$\delta^{G}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ $-2\lambda(2n-p+1)\left[(\psi\overline{\phi})_{p-1}+(\phi\overline{\psi})_{p-1}\right]$ (107)
@@
Then, we show that generalized $p$-form bilinears constructed out of generalized Killing spinors correspond to generalized KY forms when $p$ is odd and to generalized closed conformal KY forms when $p$ is even. In fact, generalized KY equation and generalized closed conformal KY equation are special cases of the generalized conformal KY equation which is written as
@@
Hence, if we have a sum of even and odd generalized $p$-form bilinears constructed out of generalized Killing spinors, then they correspond to generalized conformal KY forms.
@@ VII Conclusion
From the definition of spinor inner product for generalized spinors, we find the properties of generalized spinor bilinears for different cases. By extending the concepts of geometric Killing spinors and twistor spinors to the framework of generalized geometry, we construct the equations corresponding to generalized Killing and twistor spinors in terms of differential forms and especially we show that generalized Killing spinor equation corresponds to conformal KY equation with torsion. We extend the concept of KY forms to generalized geometry by proving that the generalized Killing equation written in terms of the generalized Lie derivative can also be described in terms of the generalized connection. Moreover, we show that bilinear forms of generalized Killing spinors with different parity correspond to generalized KY forms and closed conformal KY forms and in general they correspond to conformal KY forms for mixed parity.
As in the case of KY forms and geometric Killing spinors which can be used in the construction of extended Killing superalgebra structures, the construction of generalized KY forms and their relation with generalized Killing spinors open the way to investigate the construction of superalgebra structures generated by generalized hidden symmetries. These superalgebra structures can also be used in the description of various supergravity backgrounds. Moreover, similar constructions can be extended to the framework of exceptional generalized geometry to obtain classification results for supergravity backgrounds in various dimensions.

### v2

@@ IV Generalized Killing and twistor spinors
We now consider the action of the generalized connection on spinors. It is worth noting that while the supersymmetry parameters in supergravity backgrounds are typically sections of the spinor bundles associated with the $V_{\pm}$ splitting of the generalized metric, we here investigate the properties of the full $SO(n,n)$ spinors. These are representable as inhomogeneous differential forms on $M$. We treat these objects as the natural mathematical generalization of geometric spinors within the framework of the generalized tangent bundle $E$.
@@
where $\epsilon=\pm$ on subbundles $V_{\pm}$. In fact, the second term on the right hand side corresponds to the Clifford bracket action of $\frac{1}{2}i_{X_{a}}H$ to 1-forms $\widetilde{Y}$ and $\eta$
@@
$\alpha,\omega]_{Cl}=-2i_{X_{a}}\alpha\wedge i_{X^{a}}\omega$ (35)
and for the cases of 2-form $\frac{1}{2}i_{X_{a}}H$ and 1-forms $\widetilde{Y}$ and $\eta$, we have
@@
$i_{X_{a}}H,\mathcal{Y}]_{Cl}=\left(\begin{array}[]{cc}\widetilde{[i_{X_{a}}H,\widetilde{Y}]}_{Cl}\\ {[i_{X_{a}}H,{\eta}]}_{Cl}\end{array}\right).$ (38)
Let us consider a generalized geometry spinor $\psi$ corresponding to an inhomogeneous differential form which satisfies the Killing spinor equation in terms of the generalized connection and generalized 1-forms
@@
where we have used the identities $d^{H}=e^{a}\wedge\nabla^{H}_{X_{a}}$ and $e^{a}\wedge i_{X_{a}}\psi_{p}=p\psi_{p}$. Similarly, by applying $i_{X^{a}}$ to (48), we obtain
@@
which corresponds to the conformal Killing-Yano (CKY) equation written in terms of a connection modified by a 3-form flux $H$ corresponding to torsion [30, 31, 32, 33]. Here, $H$ corresponds to the skew-symmetric torsion and torsion 2-forms $T^{a}$ are defined as $T^{a}=\frac{1}{3}i_{X^{a}}H$. Then, a solution of the generalized geometry Killing spinor equation corresponds to a solution of the CKY equation with torsion $H$.
Similarly, let us consider the generalized geometry twistor spinor equation
@@
and the generalized Dirac operator acting on generalized geometry spinor $\psi$ corresponds to
@@
$=$ $(e_{a}\wedge+\epsilon i_{X_{a}})\nabla^{H}_{X^{a}}\psi$
$=$ $\{(e_{a}\wedge+i_{X_{a}})\nabla^{H}_{X^{a}}\psi,(e_{n+a}\wedge-i_{X_{n+a}})\nabla^{H}_{X^{n+a}}\psi\}$
$=$ $\{d^{H}\psi-\delta^{H}\psi,d^{H}\psi+\delta^{H}\psi\}$
@@
So, a solution of the generalized twistor equation corresponds to a solution of the equation (57). In particular, if the generalized geometry twistor spinor $\psi$ consists of a homogeneous degree differential form, namely it only includes the $p$-form component $\psi=\psi_{p}$, then the above equation reduces to
@@ A $p$-form $\omega$ is called a KY $p$-form, if it satisfies the following equation
because $\mathcal{G}(\mathcal{Y},\mathcal{Z})$ is a function. Since $\mathbb{D}$ is metric compatible $\mathbb{D}_{\mathcal{X}}\mathcal{G}=0$ and from (62) we have
@@
By using torsion-free condition, we can write the Courant bracket $\mathcal{X},\mathcal{Y}]_{C}$ in terms of the generalized connection $\mathbb{D}^{H}$ and we have
@@
$d<\mathcal{X},\mathcal{Y}>,\mathcal{Z}>=i_{\mathcal{Z}}d<\mathcal{X},\mathcal{Y}>=\mathbb{D}^{H}_{\mathcal{Z}}<\mathcal{X},\mathcal{Y}>=<\mathbb{D}^{H}_{\mathcal{Z}}\mathcal{X},\mathcal{Y}>+<\mathcal{X},\mathbb{D}^{H}_{\mathcal{Z}}\mathcal{Y}>$ (75)
$\mathcal{Y},d<\mathcal{X},\mathcal{Z}>>=i_{\mathcal{Y}}d<\mathcal{X},\mathcal{Z}>=\mathbb{D}^{H}_{\mathcal{Y}}<\mathcal{X},\mathcal{Z}>=<\mathbb{D}^{H}_{\mathcal{Y}}\mathcal{X},\mathcal{Z}>+<\mathcal{X},\mathbb{D}^{H}_{\mathcal{Y}}\mathcal{Z}>.$ (76)
@@
[…]
@@
[…]
@@
$(\mathbb{L}_{\mathcal{Y}}\mathcal{G})(\mathcal{X},\mathcal{X}_{A})=\frac{1}{2}\bigg(\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}}\mathcal{Y},\mathcal{X}_{A})+\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}_{A}}\mathcal{Y},\mathcal{X})\bigg)+\frac{1}{2}\bigg(\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}_{A}}\mathcal{X},\mathcal{Y})+\mathcal{G}(\mathbb{D}^{H}_{\mathcal{X}}\mathcal{X}_{A},\mathcal{Y})\bigg).$ (82)
@@
Let us choose $\mathcal{X}=\mathcal{X}_{B}$ as a basis vector which is in the same direct sum component $V_{\pm}$ with $\mathcal{X}_{A}$, and consider the admissible metric with basis $\mathcal{X}_{A}=\{X_{a}+\epsilon e^{a}\}$ and $\mathcal{E}^{A}=\{e^{a}+\epsilon X_{a}\}$. Then, we have from (20)
@@
with respect to an admissible basis $\mathcal{X}_{A}=\{X_{a}+\epsilon e^{a}\}$ which is in the same direct sum component $V_{\pm}$ with $\mathcal{K}$. So, (86) corresponds to the generalized Killing equation in terms of the generalized connection. Since $\mathcal{X}_{A}$ and $\mathcal{K}$ are in the same direct sum component $V_{\pm}$, the cross terms involving opposite direct sum components do not appear. Importantly, the equivalence of the left-hand side with the term involving the unique generalized exterior derivative $d^{H}$ ensures that the equation is well-defined, independent of the ambiguity in the specific choice of the generalized connection. As the anti-symmetric generalization of generalized Killing equation, we can write the generalized KY equation for generalized $p$-forms $\alpha$ which only includes the same direct sum components $V_{\pm}$ with $\mathcal{X}_{A}$ as
@@
where $d^{H}=\mathcal{E}^{A}\wedge\mathbb{D}^{H}_{\mathcal{X}_{A}}=d^{G}+\frac{\epsilon}{4}H\underset{1}{\wedge}$ where $d^{G}$ is defined as $d^{G}=\mathcal{E}^{A}\wedge\nabla_{\pi({\mathcal{X}_{A}})}$. Here $H\underset{1}{\wedge}=i_{X^{a}}H\wedge i_{X_{a}}$. By using (34), (87) is written explicitly as
$\mathbb{D}^{H}_{\mathcal{X}_{A}}\mathcal{\alpha}=\nabla_{X_{a}}\mathcal{\alpha}-\epsilon\frac{1}{12}[i_{X_{a}}H,\mathcal{\alpha}]_{Cl}=\frac{1}{p+1}i_{X_{a}}\left(d^{G}\mathcal{\alpha}+\epsilon\frac{1}{4}H\underset{1}{\wedge}\mathcal{\alpha}\right)$ (88)
in terms of flux terms which resembles the KY equation modified by 3-form flux $H$ corresponding to torsion. The restriction $X_{A}\in V_{\pm}$ is imposed to keep the condition compatible with the splitting defined by the generalised metric and to avoid mixing between $V_{+}$ and $V_{-}$ components in the defining equation. Importantly, this restriction should not be interpreted as implying that a metric-compatible torsion-free generalised connection is unique. Rather, the usefulness of (87) is that it admits an equivalent intrinsic formulation in terms of the induced torsionful metric connection and the $H$-twisted differential, making the dependence on an arbitrary choice of generalised Levi–Civita connection drop out at the level of the final generalised Killing–Yano condition; explicitly, (87) is equivalent to (88).
VI Relation Between Generalized KY Forms and Generalized Geometry Killing Spinors
We will investigate the relation between generalized geometry Killing spinors satisfying (39) and generalized KY forms satisfying (87). As noted previously, this construction utilizes the $SO(n,n)$ spinor representation. While distinct from the standard supersymmetry parameters in supergravity, the bilinears constructed from these generalized spinors offer a formal extension of the Killing-Yano structure to the generalized geometry framework. If we have two generalized geometry Killing spinors $\phi$ and $\psi$ satisfying (39), we can construct generalized $p$-form bilinears $(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}$ where a $p$-form bilinear is written in terms of the generalized form basis as
@@ VI Relation Between Generalized KY Forms and Generalized Geometry Killing Spinors
from the generalized geometry Killing spinor equation. Here, we use the property $\overline{\mathbb{D}_{\mathcal{X}_{A}}\psi}=\mathbb{D}_{\mathcal{X}_{A}}\overline{\psi}$. By using the definition $\overline{\psi}\phi=(\psi,\phi)$, we can write for a generalized geometry spinor $\kappa$
@@
and since $j$ is identity and $\mathcal{J}=\xi\eta$ for generalized geometry spinors, we have
@@
$d^{H}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ $2\lambda(p+1)\left[(\psi\overline{\phi})_{p+1}+(\phi\overline{\psi})_{p+1}\right]$ (102)
$\delta^{H}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ $0$ (103)
@@
which corresponds to the generalized KY equation. For $p$ is even, we have
@@
$d^{H}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ $0$ (106)
$\delta^{H}\left[(\psi\overline{\phi})_{p}+(\phi\overline{\psi})_{p}\right]$ $=$ $-2\lambda(2n-p+1)\left[(\psi\overline{\phi})_{p-1}+(\phi\overline{\psi})_{p-1}\right]$ (107)
@@
Then, we show that generalized $p$-form bilinears constructed out of generalized geometry Killing spinors correspond to generalized KY forms when $p$ is odd and to generalized closed conformal KY forms when $p$ is even. In fact, generalized KY equation and generalized closed conformal KY equation are special cases of the generalized conformal KY equation which is written as
@@
Hence, if we have a sum of even and odd generalized $p$-form bilinears constructed out of generalized geometry Killing spinors, then they correspond to generalized conformal KY forms.
@@ VII Conclusion
From the definition of spinor inner product for generalized geometry spinors, we find the properties of generalized spinor bilinears for different cases. By extending the concepts of geometric Killing spinors and twistor spinors to the framework of generalized geometry, we construct the equations corresponding to generalized geometry Killing and twistor spinors in terms of differential forms and especially we show that generalized Killing spinor equation corresponds to conformal KY equation with torsion. We extend the concept of KY forms to generalized geometry by proving that the generalized Killing equation written in terms of the generalized Lie derivative can also be described in terms of the generalized connection. Moreover, we show that bilinear forms of generalized Killing spinors with different parity correspond to generalized KY forms for the odd case and closed conformal KY forms for the even case and in general they correspond to conformal KY forms for mixed parity.
As in the case of KY forms and geometric Killing spinors which can be used in the construction of extended Killing superalgebra structures, the construction of generalized KY forms and their relation with generalized Killing spinors open the way to investigate the construction of superalgebra structures generated by generalized hidden symmetries. While the physical interpretation of these $SO(n,n)$ spinor bilinears differs from the standard bilinears of supergravity Killing spinors (which reside in the split bundle $V_{\pm}$), the mathematical parallels established here suggest that these generalized structures may provide a useful formal framework for analyzing hidden symmetries in generalized geometry.